Session: Nov 1997 (session 4)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (A.Cattani at Tue Nov 25 00:08:48 MET 1997) Success

Onsala (John Conway at Tue Nov 25 00:15:52 MET 1997) Success
No known problems. Note that no PREOBs in snap file, so no Tsys originally, added TSYS to POSTOB to get valid Tsys.

Urumqi (Li Suqin at Tue Nov 25 03:52:53 MET 1997) Success
It is fine day. H-maser to GPS offset is 2.36us. Format to GPS offset is 2.24us. No problem.

Torun (K.Borkowski at Tue Nov 25 09:12:21 MET 1997) Success

Effelsberg (dgraham at Tue Nov 25 10:03:00 MET 1997) Minor failures
Tsys measurements overflow on RCP. IFD statements inserted every 90 sec to get more frequent tsys.

Westerbork (Tony Foley at Tue Nov 25 14:17:44 MET 1997) Success
Looks OK. Phasecal switched off (for OH experiments later that night). Please let us know about the usability of the gps-fmout readings.

Noto (C.Trigilio at Tue Nov 25 17:04:58 MET 1997) Success
No problem reported.

Jodrell (Alastair Gunn at Thu Nov 27 12:41:38 MET 1997) Success
RFI in VCs 9 and 10. No other known problems.

Calibration (Fredrik T Rantakyro at Tue Dec 9 14:56:00 MET 1997) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the NOV97 directory under the name c97l4.antab Read the comments in the file to get further information.

JIVE (Lorant Sjouwerman at Wed Dec 17 16:55:34 MET 1997) Success
Network monitoring report of L-band session